OpenStudy (anonymous):

What property is shown by the equation? Column A Column B --- ABC 1. (34 + 11) + 9 = (11 + 34) + 9 --- ABC 2. 25 + (–15 + 4) = (25 + (–15)) + 4 --- ABC 3. –(–21 + 8) = 21 + (–8) --- ABC 4. –30 + (–16) = –16 + (–30) A. commutative property B. associative property C. opposite of a sum property

OpenStudy (midhun.madhu1987):

a + b = b + c Commutative Property

OpenStudy (midhun.madhu1987):

a + (b + c) = (a + b) + c Associative Property

OpenStudy (anonymous):

thanx